# Expressivity of Representation Learning on Continuous-Time Dynamic Graphs: An Information-Flow Centric Review

### v1

@@ F.1 Synthetic Datasets
To perform further experimentation and to validate our theoretical insights, we considered evaluated our considered benchmarks on a number of graph types. Since finding accessible real datasets has been a challenge, we have opted to rather use synthetic and generated datasets in which we can control the different settings and confirm our hypothesis. In the following section, we describe in details our data generation process and we give additional information to easily reproduce this process.
@@ B N $N_{B}$ $p_{i,j}$ $p_{i,i}$ $t_{gen,kc}$
For the Bi-Partite datasets, we have focused on using the real world graph TGBL-Wiki dataset, which is part of the TGB benchmark datasets (Huang et al., 2024). The dataset is a bi-partite network where Wiki pages and editors are nodes, while an edge is added when a user edits a page at a specific timestamp. For both the real and synthetic datasets, characteristics and information about the graphs utilized in the experimental results of the study are presented in Table 5.
@@
F.2 Experimental Details
@@
Finally, for our experiments on the community-based SBM dataset, we have set the number of layers to $L=1$ and a unique attention head. Similarly to the previous experiment, we have used the Adam optimizer (Kingma & Ba, 2015) with a 1e-04 learning rate and a batch size of 200. All the models were trained for $30$ epochs with the patience set to 5, and we have used a hidden dimension of $100$. All the experiments were run 5 times to reduce the effect of initialization and other randomization parameters. For CTAN, epsilon is 1.0 and gamma is 0.1, as reported in  Gravina et al. (2024).

### v2

@@ G.1 Synthetic Datasets
To perform further experimentation and to validate our theoretical insights, we evaluated our considered benchmarks on a number of graph types. Since finding accessible real datasets has been a challenge, we have opted to rather use synthetic and generated datasets in which we can control the different settings and confirm our hypothesis. In the following section, we describe in details our data generation process and we give additional information to easily reproduce this process.
@@ B N $N_{B}$ $p_{i,j}$ $p_{i,i}$ $t_{gen,kc}$
Bi-Partite Datasets. For this category of graphs, we have focused on using the real world graph TGBL-Wiki dataset, which is part of the TGB benchmark datasets (Huang et al., 2024). The dataset is a bi-partite network where Wiki pages and editors are nodes, while an edge is added when a user edits a page at a specific timestamp.
Homogeneous Graphs. We additionally experimented with homogeneous graphs, i.e. graphs where nodes and edges belong to the same type. This type of graph encompasses a wide range of applications, including social networks, transportation systems, and other domains where graph structures are uniform. For this evaluation, we used an adaptation of the real world TGBL-Coin dataset which is part of the TGB benchmarks (Huang et al., 2024) too. This dataset contains cryptocurrency transactions, where each node is an address and each edge denotes the transfer of funds from one address to another at a certain time. The results are reported in Table 7.
For both the real and synthetic datasets, characteristics and information about the graphs utilized in the experimental results of the study are presented in Table 6.
@@
Table 7: Average Test and Validation MRR ($\pm$ denotes standard deviation) of the different approaches on an adaptation of the real-world dataset TGBL-Coin. Best performance in bold.
Model Homogeneous Graph
Test Val
JODIE 0.713 $\pm$ 0.028 0.552 $\pm$ 0.131
DyRep 0.688 $\pm$ 0.007 0.697 $\pm$ 0.007
TGN 0.738 $\pm$ 0.007 0.731 $\pm$ 0.006
TGAT 0.748 $\pm$ 0.025 0.765 $\pm$ 0.023
CTAN 0.709 $\pm$ 0.003 0.754 $\pm$ 0.006
G.2 Experimental Details
@@
Finally, for our experiments on the community-based SBM dataset and homogeneous adapted TGBL-Coin dataset, we have set the number of layers to $L=1$ and a unique attention head. Similarly to the previous experiment, we have used the Adam optimizer (Kingma & Ba, 2015) with a 1e-04 learning rate and a batch size of 200. All the models were trained for $30$ epochs with the patience set to 5, and we have used a hidden dimension of $100$. All the experiments were run 5 times to reduce the effect of initialization and other randomization parameters. For CTAN, epsilon is 1.0 and gamma is 0.1, as reported in  Gravina et al. (2024).
The source code necessary to reproduce our experiments is publicly available at: https://github.com/king/ctdg-info-flow